OpenStudy (anonymous):

the perimeter of a college basketball court is 90 meters and the length is twice as long as th width. what are the length and width? please and thanks!

OpenStudy (dumbcow):

perimeter = 2w + 2L =90 2(w+L) = 90 w+L = 45 --> L = 2w w+2w = 45 3w = 45 w = 15

OpenStudy (anonymous):

thanks so much !:D